OpenStudy (anonymous):

Show that the singular point of the following function is a pole. Determine the order m of that pole and the corresponding residue B. f(z)=(1-coshz)/(z^3)

OpenStudy (anonymous):

I really wish my university had an applied complex variables class :/

OpenStudy (anonymous):

i think we can at least start this if i recall \(\cosh(z)=\cos(iz)\) if so, we can expand in a taylor series

OpenStudy (anonymous):

we can try \[\cosh(z)=\sum_{k=0}^{\infty}\frac{z^{2k}}{(2k)!}\]

OpenStudy (anonymous):

this should give \[1-\cosh(z)=-\sum_{k=1}^{\infty}\frac{z^{2k}}{(2k)!}\]

OpenStudy (anonymous):

yeah got it

OpenStudy (anonymous):

Which would then give: \[-\sum_{k=1}^{\infty}\frac{z^{2k-3}}{(2k)!}\] right?

OpenStudy (anonymous):

divide by \(z^3\) first term is \(\frac{1}{2z}\) and all the others have \(z\) in the numerator

OpenStudy (anonymous):

@malevolence19 yes, that is what i get too. or shift the index either way.

OpenStudy (anonymous):

So the the residue is just 1/2?

OpenStudy (anonymous):

so b1 comes out to be -1/2.yeah?

OpenStudy (anonymous):

yes if my memory serves me in this context the residue is the coefficient of the \(z^{-1}\) term

OpenStudy (anonymous):

^right

OpenStudy (anonymous):

Yeah, -1/2 i forgot about that in front.

OpenStudy (anonymous):

i would not place much money on this, because it has been a while, but i am pretty sure it is correct

OpenStudy (anonymous):

and pole is??i don't have any idea what the heck is it

OpenStudy (anonymous):

It would just be zero right?

OpenStudy (anonymous):

It it was like: \[\frac{B}{(z-\lambda)}\] then the pole would be lambda? I think, but I'm going off very little knowledge.

OpenStudy (anonymous):

lambda is a singular point..not sure about pole

OpenStudy (anonymous):

yes, the pole is at 0, because that is where the denominator is undefined. it is a simple pole because it is a zero of order one

OpenStudy (anonymous):

^the denominator becomes undefined at singular points...so every singular point is a pole??

OpenStudy (anonymous):

i think its pole is 1

OpenStudy (anonymous):

a "pole" is a kind of singular point think of a simple example from pre calc class \[\frac{x^2-4}{x-2}\] vs \[\frac{x^2}{x-2}\] in the first example, 2

OpenStudy (anonymous):

this thing implies m=1 at b1

OpenStudy (anonymous):

in both examples, 2 is a singular point because the function is not defined at 2 but in the first example, 2 is a removable singularity, because you can remove it

OpenStudy (anonymous):

in the second example, 2 is a pole, in pre calc a "vertical asypmtote"

OpenStudy (anonymous):

however it is spelled

OpenStudy (anonymous):

isolated singular point.what is it??

OpenStudy (anonymous):

ok hold the phone

OpenStudy (anonymous):

what you posted was about the order of the pole, not the pole itself

OpenStudy (anonymous):

in your case above, we had \(\frac{1}{2z}\) plus stuff with positive exponents, means you have a pole of ORDER 1, not a pole at \(z=1\)

OpenStudy (anonymous):

here is a nice succinct explanation (very short) with example http://www.solitaryroad.com/c612.html

OpenStudy (anonymous):

implies order of pole=k if bk is the co-efficient of z^-1 term.right?

OpenStudy (anonymous):

the order of the pole is the multplicty of the zero on the denominator, to put it in pre calculus terms, example \[\frac{1}{(z-5)^3}\] has a pole of order 3 at \(z=5\)

OpenStudy (anonymous):

\[\frac{1}{z-2}\] has a pole of order 1 at \(z=2\) a pole of order one is called a simple pole

OpenStudy (anonymous):

oh *shouts got it*

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

thanks bro

OpenStudy (anonymous):

yw